OpenStudy (anonymous):

x + y = 4 x - 2y = -5 If the first equation is multiplied by 2 and then the equations are added, the result is _____. answers 3x = 13 3x = 3 x = 3

jhonyy9 (jhonyy9):

2x+2y=8 x-2y=-5 ---------- 3x 0 =3 3x=3 x=3/3 x=1

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

x=1 y=3